OpenStudy (anonymous):

Please help my last question. Find x for the following proportion, 2x-3/3=2/3 Answers, 1/4 3 5/2 6

OpenStudy (anonymous):

asking for help for a quiz is cheating on a side note, notice how the denominators of the 2 fractions are the same

OpenStudy (anonymous):

yes i do

OpenStudy (whpalmer4):

\[\frac{2x-3}{3} = \frac{2}{3}\] You can multiply both sides of the equation by 3 to eliminate the fraction. Can you solve the resulting equation?